OpenStudy (anonymous):

find the laplace transform of sin2tcos2t?

OpenStudy (ksaimouli):

u mean derivative

OpenStudy (anonymous):

no it is the laplace transform.

OpenStudy (anonymous):

master level math

OpenStudy (anonymous):

university

OpenStudy (accessdenied):

I am not particularly familiar with Laplace Transforms, but I am curious if it would be helpful to simplify sin(2t)cos(2t) using a trig identity: \(\sin(2x) = 2 \sin x \cos x\) \(\sin 2t \cos 2t = \frac{1}{2}\left( 2 \sin 2t \cos 2t \right) = \frac{1}{2} \sin (4t) \) At least, it tends to be easier to play with a single sine than a product of two trig functions. :)

OpenStudy (anonymous):

So , 1/2 w/s^2 +w^2 where w = 4 -> 2/(s^2+16)